OpenStudy (anonymous):

Given C = 118, B = 50, and c = 15, use the law of sine to solve the triangle for the value of a... would I use the formula a/ sin A = b/ sin B?

hero (hero):

sin C/c = sin B/b

OpenStudy (anonymous):

A^2=B^2+C^2-2ab cos a

OpenStudy (anonymous):

hero I used that formula earlier and I got b = 6.355932203

OpenStudy (anonymous):

i know i have to plug this in to a/ sin A = b/ sin B but I keep getting the wrong answer

hero (hero):

118 + 50 + A = 180 solve for A

OpenStudy (anonymous):

\[\frac{b}{\sin(50)}=\frac{15}{\sin(118)}\] \[b=\frac{15\sin(50)}{\sin(118)}\] make sure your calculator is in "degree" mode since you are working with degrees

OpenStudy (anonymous):

i get something pretty close to 13 http://www.wolframalpha.com/input/?i=15sin%2850%29%2Fsin%28118%29

hero (hero):

I'm used to the sin being on the top, not the bottom. Guess it really doesn't matter

OpenStudy (anonymous):

okay I worked it all out and i got a = 3.12?? but I know that's not right..

hero (hero):

Did you find angle A?

OpenStudy (anonymous):

yes. C is 118 and B is 50. I added those together and subtract it from 180 and it gave me 12

hero (hero):

Angle A is so small. I expect its side to be small as well.

OpenStudy (anonymous):

yes. so I have a/12 = 13/50...I multiply 12 and 13 and got 156. divide that by 50 i get 3.12..but that's not the right answer

hero (hero):

Must be the ambiguous case

OpenStudy (anonymous):

my choices are 17.29, 4.76, 4.07, 2.31, 3.53...obviously its not going to be 17.29

OpenStudy (anonymous):

so that can be ruled out

OpenStudy (anonymous):

this line so I have a/12 = 13/50 is not correct

OpenStudy (anonymous):

what was my mistake?

OpenStudy (anonymous):

we found \(b\) right? \(b=13\)

OpenStudy (anonymous):

yes and the formula is a/ sin A = b/ sin B

OpenStudy (anonymous):

ok so \(\frac{a}{\sin(15)}=\frac{13}{\sin(50)}\) making \(a=\frac{13\sin(50)}{\sin(15)}\)

OpenStudy (anonymous):

okay, so im confused as of what to do here. divide?

OpenStudy (anonymous):

first of all you are supposed to ignore me because that equation was wrong

OpenStudy (anonymous):

\(\frac{a}{\sin(15)}=\frac{13}{\sin(50)}\) so \(a=\frac{13\sin(15)}{\sin(50)}\)

OpenStudy (anonymous):

that is better

OpenStudy (anonymous):

i got 3.9 now

OpenStudy (anonymous):

i get \(4.4\) rounded http://www.wolframalpha.com/input/?i=13sin%2815%29%2Fsin%2850%29

OpenStudy (anonymous):

are you using a calculator? easier to put it in wolfram in one step

OpenStudy (anonymous):

yes i am..and neither of the answers we are getting is one of my choices..

OpenStudy (anonymous):

what are your choices for \(b\) ?

OpenStudy (anonymous):

i only have choices for a since that is what i am looking for. when i solve for b i got 6.38..idk how you are getting 13

OpenStudy (anonymous):

ok lets go slow, see if we made any mistakes you have \(C=118^o, c=15, A=50^o\) is that right?

OpenStudy (anonymous):

no B = 50, not A

hartnn (hartnn):

sat, only 1, A = 12 so u have sin 12, not sin 15

OpenStudy (anonymous):

oh ok \(B=50^o\) and so \(A=12^o\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@hartnn so it is!!!

hartnn (hartnn):

yes, and now u get an answer which is in choices..

OpenStudy (anonymous):

then we know \(\frac{a}{\sin(12)}=\frac{15}{\sin(118)}\) and so \[a=\frac{15\sin(12)}{\sin(118)}\]

OpenStudy (anonymous):

i get \(3.53\) http://www.wolframalpha.com/input/?i=15sin%2812%29%2Fsin%28118%29

hartnn (hartnn):

yup, thats it.

OpenStudy (anonymous):

I got the same too, finally! thank you!

OpenStudy (anonymous):

imagine!

hero (hero):

Well, at least the right answer was found